OpenStudy (anonymous):

On a military (24-hour) clock, what time is it 3122 hours after 16 hundred hours?

OpenStudy (shane_b):

So you're adding 3122 hours to 1600 which = 4722. 4722 % 24 = 0.75 which is 3/4 of a day...which is 1800 in military time.

OpenStudy (anonymous):

Aren't you adding 3122 to 16 = 3138/24 = 130 with remainder 18, so it is 1800 hours that way?

OpenStudy (shane_b):

Yes. I suppose I goofed that up when I typed it out. It should have been (3,122 + 16) % 24=0.75 = 1800 hours (6pm). fixed typo :P